HUBBLE SPACE TELESCOPE DAILY REPORT #982 PERIOD COVERED: 0800 EST Daily Status Report as of 272/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed GHRS/2 3859 (Nighttime Measurements of Nitric Oxide - Continuation of 2544) The High Resolution Spectrograph was used to make observations of the bright B class star HD037128. It is hoped that these observations will help learn more about the diffusion processes in the Earth's upper atmosphere by studying the ultraviolet absorption bands produced by nitric oxide along the line of sight of the target star. The observations was reported to be rather weak. The OSS said that they suspect that the target was not completely in the aperture. 1.2 Completed GHRS/2 4124 (GHRS Sensitivity Monitor - Cycle 2) The High Resolution Spectrograph was used to make a series of observations using standard star MC-Col to look for changes in the instrument's sensitivity. The onboard target acquisition was successful. All following 64 observations were reported to look normal. 1.3 Completed PC 5014 (WF/PC Cycle 3 Calibration: PC Flats Narrow. 2) The Planetary Camera was used to make standard Earth calibrations images for flat fielding of PC observations. Medium and narrow filters were used, with and without neutral density filters. The observations were executed as scheduled. Four observations were received and displayed. One was reported to be saturated, and three were normal. 1.4 Uplinked Updated Ephemeris Tables, Per MOSES 2871 1.5 Completed PC 5072 (WF/PC Cycle 3 Calibration: PC Flats Wide. 1) Both the Wide Field Camera and the Planetary camera were used to make Earth Flat calibration exposures using the wide and medium band filters. Six observations were received and displayed in the OSS. Three were reported to be normal, one was underexposed, and two were partially saturated. 1.6 Completed FGS 4893 (Orbital Motions of Close Binary Stars - Cycle 3 High) Fine Guidance Sensor #3 was used to make astrometric observations of the binary star system HD195987 to obtain its orbital elements. These observations are important since most of the estimates of stellar mass are made by observing elements of these systems. The astrometry was monitored and no problems were seen in the telemetry. 1.7 Completed WFC 4848 (GTO Low Latitude Parallel Program - Cycle 3 High) The Wide Field Camera was used to make two high latitude medium-deep sky survey observations taken in parallel with the prime science observations. The observations were executed as scheduled, and no problems were reported. 1.8 Completed PC 5017 (Internal Velocity Distribution in Globular Clusters - Part three - Cycle 3) The Planetary Camera was used to make observations of Globular Cluster NGC-6205. This set of observations will be used along with observations from proposal 1007, to study the proper motion of this cluster. Observations were made at different distances from the core center out to ten core radii. The observations were monitored in the OSS, and the data quality was assessed. No problems were reported. 1.9 Completed WFC/PC 5026 (WF/PC Cycle 3 Calibration: Rapid Internal Monitor) The Wide Field Camera and the Planetary Camera were used to make internal flat calibration images. These images will be used to look for measles and daisies (i.e., different forms of contamination), and to monitor scattered light. The observations were executed as scheduled, and no problems were reported. 1.10 Completed FOS/BL 4260 (Extreme Stellar Winds and Post - Main - Sequence Evolution) The Faint Object Spectrograph was used to make observations of star BE335. This star is believed to be intermediate between a class O and WR star and to have strong winds and high mass loss rate (i.e., much higher than normal O class stars). The observations were executed as scheduled, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 3 Successful : 3 Scheduled Reacquisitions : 8 Successful : 8 2.2 FHST Updates: Scheduled : 10 Successful : 10 2.3 Significant HSTARs: None 2.4 Operations Note: None The JIS #3 Simulation (Deployment) originally scheduled for yesterday (Tuesday Sept. 28, 1993), was canceled and was rescheduled for today (Wednesday Sept. 29,1993). It started at 1300Z, and should last for ten hours. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete Four Sets of WFC 5075 (WF/PC Cycle 3 Calibration: (WFC Flats Wide. 1) 3.2 Complete Three Sets of WFC 4381 (HST Medium - Deep Survey - Cycle 3) 3.3 Complete FOS/BL 3441 (Post Asymptotic Giant Branch Evolution in the Magellanic Clouds) 3.4 Complete PC 5027 (WF/PC Cycle 3 Calibration: Rapid Internal Monitor - Part 2) 3.5 Complete Two Sets of WFC 5013 (WF/PC Cycle 3 Calibration: WFC Flats Narrow. 2) 3.6 Complete FGS 4938 (Parallaxes of Hyades Cluster Members - Cycle 3 High) 3.7 Complete WFC 4848 (GTO Low Latitude Program - Cycle 3 High) 3.8 Complete FOC/96 3881 (Crane, Augmentation, Galaxies & Clusters) 3.9 Complete WFC 4849 (GTO High Latitude Parallel Program - Cycle 3 High) 3.10 Complete FGS 4893 (Orbital Motions of Close Binary Stars - Cycle 3 High) 3.11 Complete PC 4764 (Investigation of the Dynamics of Binary Asteroids) 3.12 Complete PC 5014 (WF/PC Cycle 3 Calibration: PC Flats Narrow. 2) 3.14 Complete FOC/96 3504 (High - Spatial - Resolution Imaging and Spectroscopy of AGN: Cycle 2) 3.15 Complete WFC 5201 (GTO High Latitude Parallel Program - Cycle 3 High - Continued) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: The VEST/SEER Internal simulation #4 Repeat 10 (EVA 3) which was originally scheduled for today (Wednesday, Sept. 29, 1993), has be rescheduled for tomorrow (Thursday, Sept. 30, 1993) at 1 PM until 9 PM. /DMH -END- .